Home‎ > ‎Reports‎ > ‎Crew101‎ > ‎

day11


Crew 101 Log Book for March 22, 2011

Photos of the day





Commander's Check-In Report

Tom Caillouet Reporting

Crew Health: Quite good

EVAs and Times:

EVA 13: 1148 - 1255

EVA 14: 1353 - 1416

EVA 15: 1416 - 1608

Number of Reports:8 (Commander, Engineering, Journalist, Chef, 3 EVA, Science)

Field Results:

EVA 13 - Successfully repaired radio repeater after wind damage.

EVA 14 - Showed the beginning of Radio Ridge road to Cody in preparation of EVA 15

EVA 15 - Testing of radio range on Radio Ridge trail. Forced to cut short early due to inclement weather.

Plans for Tomorrow:Morning EVA to retry testing of radio range on Radio Ridge trail. Afternoon exploratory collection EVA. Continuing biology science in-Hab.


Commander's Report

Tom Caillouet Reporting

The crew almost seems to be winding down. We can feel that our time here is limited. Biology samples have been completely gathered for days. We could get more geology samples, but the investigator back at Tech is already pleased with the current haul. Radio coverage would have been completely tested today, if it weren't for uncooperative weather (rain and snow cut that EVA short), but it was still as very successful EVA as the radios demonstrated a marked improvement with a longer antenna. I'm pleased that we've reached this point with several days left in the mission. We will be able to use the last couple days to do more exploratory EVAs and use the areohashing algorithm, which has been more or less shelved while we finalize the other experiments. Still, it is disappointing that we have only a few more days in sim, if only because we can't escape the work that expected of us on Earth for much longer. No point in taking our eyes too far off of our current activities. We've got several EVAs to go and a lot of science and writing to do over the next couple days.


Engineering Report

Cody Hall Reporting

Generator/Electricity:

• Kitty (Diesel Generator) (Used? Yes/No): Yes

• Honey (Gas Generator) (Used? Yes/No): No

• Battery Charge Level:

o V everstart

o V armor plate

o V green

Notes/Comments (include how many generator runs, approximately when, and any times you needed to turn the generator on/off):

1 generator run

Fuel Status:

• Diesel (Full/0.75/0.5/0.25/Empty): 0.75

• Propane Tank (Full/0.75/0.5/0.25/Empty): 0.25

• Gasoline Tank (Full/0.75/0.5/0.25/Empty): 0.5

Notes/Comments:

Oil:

• Kitty Oil Quantity (# of Quarts in storage): Changed/replaced by DG

• Honey Oil Quantity (# of Quarts in storage): Leave it to DG

• ATV Oil Quantity (# of Quarts in storage): 6 quarts on the engineering work bench

• Car Oil Quantity (# of Quarts in storage): Leave it to DG

Notes/Comments:

Water Status:

• Outside Potable Water Tank Level (inches from bottom): 9

• Trailer Potable Water Tank Level (inches from bottom): full

• Hab Potable Water Tank Level (inches from bottom): 12

• Potable Water Meter Reading: 437804

Notes/Comments :


GreenHab:

• Condition of plants in Tank 1: unchanged

• Condition of plants in Tank 2: unchanged

• Condition of plants in Tank 3: unchanged

• Greenhab notes/ comments:

Observatory:

• Telescope Used? (Yes/no) No

Observatory Notes:


Transportation:

Hab Car (NOT RENTAL)

• Used (yes/no): No

• Oil Checked (yes/no): No

• Tire Status: Not checked

• Odometer at end of day: 123688.4

• Notes/Comments on Hab Car:

ATV 1 (Opportunity):

• Used: yes

• Oil Checked: no

• Fuel Consumed: replaced

• Tire Status: ok

• Comments:


ATV 2 (Spirit):

• Used: yes

• Oil Checked: no

• Fuel Consumed: replaced

• Tire Status: ok

• Comments:

ATV 3 (Viking I):

• Used: no

• Oil Checked: no

• Fuel Consumed: none

• Tire Status: ok

• Comments: think battery may be dead; could not start


ATV 4 (Viking II):

• Used: no

• Oil Checked: no

• Fuel Consumed: none

• Tire Status: ok

• Comments:

ATV Comments:

Heating and Ventilation:

• Thermostat setting upstairs (Farenheit): 68

• Actual temperature upstairs (Farenheit): 70

• Thermostat setting downstairs (Farenheit): 67

• Actual temperature downstairs (Farenheit): 65

Notes/Comments:

Computers/Networking Infrastructure: working

Notes/comments on internet/computers:


General Engineering

Notes/Comments: found rain collector from hobo outside. It probably fell off during high winds yesterday. Did not attempt to reattach it.


Science Report

Christina Graves Reporting

Objectives:

Science objectives for the day is to complete DNA isolation from soil samples collected thus far (8 in al) using the E.Z.N.A. Soil DNA kit (Omega). Also prepare liquid cultures for gram staining.

Materials and Equipment:

1. water bath set to 70C

2. centrifuge

3. 1.5 mL centrifuge tubes

4. 15mL centrifuge tubes

5. 95% ethanol for sterilization

6. ice bath

7. 91% isopropanol

8. LB mix

9. dIH2O

10. pressure cooker

11. sterile inoculation loops


Preparation and Procedure:

The water bath was set to 70C. Prepare an ice bath using a styrofoam bucket and ice.

Each sample was transferred to sterile 1.5mL eppendorf tubes, and were centrifuged at 3000rpm for 3 min (Eppendorf Minispin plus centrifuge). The supernatant was then transferred to a new 1.5mL tube, and 270uL of Buffer P2 was added. The samples were then incubated on ice for 5 min, and centrifuged at full speed (13000 rpm) for 5 minutes [note that the kit protocol calls for a 4C microcentrifuge, which was unavailable, and the samples were centrifuged at RT]. This supernatant was transferred to a new 1.5mL tube, and 0.7 volumes of isopropanol (91%) was added. The sample was then incubated at 0C for 1 hour [note: protocol calls for -20C]. Samples were then centrifuged at full speed for 10 minutes [note: protocol calls for 4C centrifuge]. Cells were then allowed to incubate on ice again for 1 hour to increase DNA concentration.

Plates were also removed from the incubator. Successful bacterial colony growth for samples for all samples except sample 1.

LB media was prepared using 500mL of dIH2O, and 13g of LB media. Lid was loosely capped and placed in a pressure cooker with about 1in. of water in the bottom. The solution was heated in the pressure cooker for 45 minutes to ensure sterilization [note: the pressure cooker is missing the top; I'm not sure if that will affect sterilization technique or not].

When cooled, 1000uL of LB media was added to a 1.5mL centrifuge tube, and colonies were picked with sterile inoculation loops. Tubes were then placed in the incubator at 37C.

Results: As mentioned, successful growth was seen from all samples except sample 1. Sample 8, in particular, grew extremely well.


EVA 13 Report

Jon Kosh Reporting

EVA Time: 1148-1255

EVA Crewmembers: Jon KF7GLD, Mitesh KK4AQI

Base Communications: Tommy KJ4KAC

EVA Procedure: High winds from the weekend (gusts of up to 48mph) have brought the repeater antenna down. The crew will depart the Hab with a shovel and hacksaw to re-raise the repeater antenna.

EVA Results: Repeater antenna successfully raised. It was found that, as anticipated, the fiberglass mast holding the antenna had snapped near the base of the repeater setup, just above where the mast was buried. The EVA team sawed off the bottom portion of the mast and dug a deeper hole to hold the mast. The slightly shorter mast was inserted and raised to its full height.

Lessons Learned: Our pelican box holding our repeater equipment is sufficiently weighed down by the two 12V batteries inside, however, wind speeds in the mid 40mph range will bend our fully erect repeater to critical angles.


EVA 14 Report

Tom Caillouet Reporting

EVA Time:1353-1516

EVA Crewmembers: Tommy, Cody

Base Communications: Jon KF7GLD

EVA Procedure:The purpose of this EVA is to prepare for the later EVA along the Radio Ridge trail. Tommy will lead Cody to show the beginning of this trail. The crew will then retrieve UTM coordinates of the trailhead.

EVA Results: Cody was successfully shown the beginning of the trail. UTM coordinates found were: 12N 517812 4254502. The crew was also

surprised to learn that the new radio improvements now made two way

communication from the road with the Hab available for the length of Cow Dung Road to the trail separation.

Lessons Learned: None


EVA 15 Report

Cody Hall Reporting

EVA Crewmembers: Cody Hall, KJ4RLF;Christine Redmond, KD8PPY

Base Comms:Tommy Caillouet, KJ4KAC

EVA Time:1416-1608

Mode of Transportation: ATV

EVA Procedure:EVA team would travel north along Cow Dung Road, stopping every few minutes to give GPS coordinates and gage ham radio signal strength for newly raised repeater antenna. Once team reaches Brahe Highway, they would head west. Team would stop every few minutes to give GPS location, and base would gage signal strength. Once team reached what appeared to be Copernicus Highway, team would turn south and repeat GPS and signal strength tests. After many check-ins, team would turn around and return to base.

EVA Results:Signal strength was confirmed to be improved due to readings on Cow Dung Road. There is no signal strength to transmit or receive on Brahe Highway, or Copernicus Highway until radios are nearly due west of the Hab. Two GPS locations were taken, but only radios transmitting on high power could transmit to base. After these two readings, team noticed that clouds were getting dark overhead and decided to head back instead of risking getting caught in rain or snow.

Lessons Learned:The roads off of Cow Dung Highway are much rougher, and teams will want to drive much more carefully. Crews should probably check weather reports before planning long and far away EVAs.


Journalist Report

Christine Redmond Reporting

This morning started with a repair EVA. During yesterday’s wind storm our repeater antenna snapped at the base cutting communication for our EVAs. Mitesh and Jon suited up and climbed radio ridge to assess and repair the damage. Luckily we only lost about a foot and a half from our initial antenna length.

This morning Cody and I headed out on EVA .The EVA  was quite an adventure! We were gone for approximately two hours and in that time we traveled along a path west off of Cow Dung trail testing radio coverage in these regions previously unexplored by crew 101. These trails are much more rugged than the nearby trails making this an extremely fun EVA to maneuver. We had to cut our EVA short when dark clouds appeared over head and it started snowing!

Today I really felt like an astronaut when I added one more addition to my space suit, a heart rate monitor. Mitesh is heading a study to relate heart rate during different activities to oxygen consumption. He is hoping by studying this relationship he can determine how much oxygen would be needed to support human life on Mars with relationship to different activities that would be performed on Mars. Christina had a busy day analyzing our biology samples. Raj has been enjoying today’s beautiful Martian view. Tommy has been working on our final report.

I am happy to report an adventurous and productive day. We are all looking forward to another night on the red planet!